# High-Performance Computational Magnetohydrodynamics with Python

### v1

@@ Abstract
We present the AGATE simulation code, a Python-based framework developed primarily for solving the magnetohydrodynamics (MHD) equations while maintaining adaptability to other equation sets. The code employs a modular, object-oriented architecture that separates interface specifications from numerical implementations, allowing users to customize numerical methods and physics models. Built on a Godunov-type finite-volume scheme, AGATE currently supports the ideal, Hall, and Chew-Goldberger-Low (CGL) MHD equations, with multiple acceleration options ranging from Numpy to GPU-enabled computation via NVIDIA CUDA. Performance testing demonstrates that our GPU implementations achieve 40-60x speedups over CPU versions. Comprehensive validation through established benchmarks confirms accurate reproduction of both linear and nonlinear phenomena across different MHD regimes. This combination of modularity, performance, and extensibility makes AGATE suitable for multiple applications: from rapid prototyping to production simulations, and from numerical algorithm development to physics education.
@@
Multiple acceleration options support both CPU and GPU computation, with GPU implementations achieving 40-60x speedups over CPU versions
@@ 4.7 Timings
We conducted performance benchmarks to evaluate AGATE’s various acceleration options. Using the Orszag-Tang vortex test case (Section 4.2), we simulated ideal MHD with divergence cleaning to time $t=0.5$ across grid resolutions ranging from $32^{2}$ to $2048^{2}$. All simulations used default algorithm settings and were executed on a system featuring an AMD EPYC CPU (2.654 MHz; 2 sockets; 64 cores/socket; 2 threads/core; 256 threads total) and an NVIDIA A100 GPU. For the Numba acceleration option, preliminary testing on a $256^{2}$ grid identified optimal performance with our system CPU using 16 threads. Comparative timing results are presented in Table 2.
@@
n None Numpy Numba Cupy Kernel
$32^{2}$ 42.4 0.304 10.99 1.01 0.515
$64^{2}$ 341.7 1.52 11.17 1.13 0.779
$128^{2}$ 2714 13.76 14.41 2.08 1.272
$256^{2}$ N/A 105.0 34.65 4.02 2.272
$512^{2}$ N/A 790.8 171.5 8.12 4.525
$1024^{2}$ N/A 6614 1135 28.76 19.5
$2048^{2}$ N/A N/A N/A 216.9 133.7
@@
Table 2: Performance comparison of AGATE acceleration modes for ideal MHD Orszag-Tang vortex simulations (with divergence cleaning) to $t=0.5$. Timings are reported in seconds. CPU-based modes (None, Numpy, Numba) were executed on an AMD EPYC CPU with Numba utilizing 16 threads. GPU-based modes (Cupy, Kernel) employed an NVIDIA A100 GPU. Note: Numba requires a one-time 10-second compilation overhead for each run. ”N/A” indicates that the benchmark was not run due to prohibitively long execution times.
@@
Our performance testing reveals that while a pure Python implementation is impractical for high-performance computing, using acceleration libraries result in substantial improvements. Numpy alone provides a 200-fold speedup over pure Python for a $128^{2}$ grid. Numba, despite requiring a 10-second initial JIT compilation overhead, achieves 3-5 times better performance than Numpy for sufficiently large problems. The GPU-based options demonstrate the highest performance, with speedups of 40-60 times over Numba and 230-340 times over Numpy for our large-scale benchmark. These results confirm that our Python-based implementation can effectively support high-performance computing applications.
@@ 5 Conclusion
Performance testing reveals that while pure Python (acceleration option: None) is impractical for most uses, the Numpy and Numba CPU options provide viable speedups for small simulations. GPU acceleration through Cupy and Kernel yields further performance gains, achieving speedups of 40-60x over Numba for large problem sizes. These results demonstrate that Python-based scientific computing can achieve competitive performance when properly optimized.

### v2

@@ Abstract
We present the AGATE simulation code, a Python-based framework developed primarily for solving the magnetohydrodynamics (MHD) equations while maintaining adaptability to other equation sets. The code employs a modular, object-oriented architecture that separates interface specifications from numerical implementations, allowing users to customize numerical methods and physics models. Built on a Godunov-type finite-volume scheme, AGATE currently supports the ideal, Hall, and Chew-Goldberger-Low (CGL) MHD equations, with multiple acceleration options ranging from Numpy to GPU-enabled computation via NVIDIA CUDA. Performance testing demonstrates that our GPU implementations achieve $>$40x speedups over CPU versions. Comprehensive validation through established benchmarks confirms accurate reproduction of both linear and nonlinear phenomena across different MHD regimes. This combination of modularity, performance, and extensibility makes AGATE suitable for multiple applications: from rapid prototyping to production simulations, and from numerical algorithm development to physics education.
@@
Multiple acceleration options support both CPU and GPU computation, with GPU implementations achieving ¿40x speedups over CPU versions
@@ 4.7 Timings
We conducted performance benchmarks to evaluate AGATE’s various acceleration options, noting that our primary focus is to use the GPU options for high-performance computing applications. Using the Orszag-Tang vortex test case (Section 4.2), we simulated ideal MHD with divergence cleaning to time $t=0.5$ across grid resolutions ranging from $32^{2}$ to $2048^{2}$. All simulations used default algorithm settings and were executed on a system featuring an AMD EPYC CPU (2.654 MHz; 2 sockets; 64 cores/socket; 2 threads/core; 256 threads total) and an NVIDIA A100 GPU. For the Numba acceleration option, preliminary testing on a $256^{2}$ grid identified optimal performance with our system CPU using 16 threads. Comparative timing results are presented in Table 2.
@@
n None Numpy Numba Cupy Kernel KernX3 Ath++ AthOMP ChollaX3
$32^{2}$ 28.8 0.28 13.3 0.73 0.27 0.33 0.048 0.065 0.76
$64^{2}$ 231 1.46 13.6 1.37 0.450 0.49 0.362 0.204 0.84
$128^{2}$ 1802 17.6 16.4 2.63 0.82 0.98 2.74 0.985 1.17
$256^{2}$ N/A 122 49.5 5.17 1.55 2.66 22.1 4.26 3.16
$512^{2}$ N/A 895 247 10.3 3.04 19.0 177.0 24.4 18.0
$1024^{2}$ N/A 7477 1320 20.3 12.0 147 1396 214 132
$2048^{2}$ N/A N/A N/A 93.4 87.1 1172 N/A 1003 1021
@@
Table 2: Performance comparison of AGATE acceleration modes for ideal MHD Orszag-Tang vortex simulations (with divergence cleaning) to $t=0.5$. Timings are reported in seconds. CPU-based modes (None, Numpy, Numba) were executed on an AMD EPYC CPU with Numba utilizing 16 threads. GPU-based modes (Cupy, Kernel) employed an NVIDIA A100 GPU. KernX3 is performed on a $N^{2}x3$ grid in order to match ChollaX3, which can only operate with a 3D grid. Ath++ is Athena++ run on a single node, while AthOMP utilizes OpenMP with parameter num_threads set to 16. Note: Numba requires a one-time 13-second compilation overhead for each run. ”N/A” indicates that the benchmark was not run due to prohibitively long execution times.
@@
Our performance testing reveals that while a pure Python implementation is impractical for high-performance computing, using acceleration libraries result in substantial improvements. Numpy alone provides about a 100-fold speedup over pure Python for a $128^{2}$ grid. Numba, despite requiring a 13-second initial JIT compilation overhead, achieves 2-5 times better performance than Numpy for sufficiently large problems. The GPU-based options demonstrate the highest performance, with speedups of $>$40 times over Numba and $>$300 times over Numpy for the larger benchmarks.
We compare our timing results with established HPC codes in the literature: the CPU-based Athena++ [59, 60] and the GPU-based Cholla [61, 62]. Using the same Orszag-Tang benchmark across identical problem sizes, we followed the implementation instructions provided in each code’s repository. Athena++ supports a OpenMP implementation; for this mode, we set the size of the individual MeshBlocks such that the parameter num_threads was equal to 16 for all problem sizes. Since Cholla requires a grid with at least three dimensions, we used an NxNx3 grid for the benchmark. For a more equitable comparison with AGATE, we also executed our Kernel benchmark with a similar 3D grid configuration (KernelX3).
An advantage of our Python-based approach over the C++ implementations of Athena++ and Cholla is the virtual elimination of compilation time. AGATE’s only preprocessing overhead resulted from JIT compilation, which took approximately 13 seconds in our testing environment for the Numba option. AGATE’s GPU options had minimal (<<much-less-than$<$ < 1 second) overheads, and the other CPU options had zero overhead. This compares favorably to Athena++ and Cholla, which required 81 and 119 seconds to compile, respectively.
Our results clearly show that Athena++ is considerably more optimized for CPU-based HPC operations than AGATE, which is expected given our focus on GPU implementations. Nevertheless, AGATE’s minimal compilation requirements and seamless integration with Python workflows make it particularly suitable for rapid deployment, especially for students in learning environments or when prototyping with smaller problem sizes.
For GPU performance, we found that Cholla executes approximately 10-15% faster than AGATE at the largest problem sizes, while performance is comparable at smaller scales. We consider this modest speed difference an acceptable trade-off for the advantages of a Python-based implementation. Overall, our results demonstrate that AGATE effectively supports high-performance computing applications when leveraging GPU acceleration.
@@ 5 Conclusion
Performance testing reveals that while pure Python (acceleration option: None) is impractical for most uses, the Numpy and Numba CPU options provide viable speedups for small simulations. GPU acceleration through Cupy and Kernel yields further performance gains, achieving speedups of $>$40x over Numba for large problem sizes. These results demonstrate that Python-based scientific computing can achieve competitive performance when properly optimized.